OpenStudy (anonymous):

can someone PLEASE tell me what the derivative of: x(12-2x)^2 is?????

OpenStudy (mimi_x3):

You can use the product rule to differentiate it; or expand it out. Have you tried it?

OpenStudy (anonymous):

the easiest way to do this is to simply expand, and do each part seperately

OpenStudy (anonymous):

x(144+4x^2-48x) = 144x + 4x^3 - 48x^2 Derivative: 144 + 12x^2 - 96x

OpenStudy (anonymous):

I have tried everything. I think the answer in the book is wrong but I wanted to make sure.

OpenStudy (anonymous):

GT HOW did you get that???

OpenStudy (anonymous):

You understand how I got this one, right? 144x + 4x^3 - 48x^2 Now, derivative of that is: d/dx 144x = 144 d/dx 4x^3 = 4*3*x^2 d/dx -48x^2 = -48*2*x

OpenStudy (anonymous):

I think I had that before but I erased it.

OpenStudy (anonymous):

So, finally, you get: 144 + 12x^2 - 96x

OpenStudy (anonymous):

Yes thank you!!!! At first I thought I had to use the chain rule but I guess I was over thinking it.

OpenStudy (anonymous):

Well....you could. But, this one is simple enough to expand out.

OpenStudy (anonymous):

I wasn't getting it with the chain rule. It's been about 15 years since I have had calculus. I'm rusty! Thanks though!

OpenStudy (anonymous):

No worries.

OpenStudy (anonymous):

No, GT. Chain rule is for the composition of two functions f(g(x)). Product rule is for the product of two functions f(x)*g(x).

OpenStudy (anonymous):

Sorry.....confused the two. yeah...I meant product rule. :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

There....removed the error.